OpenStudy (anonymous):

Simplify by removing factors of 1. p^2-4 / (p+2)^2

hero (hero):

Ah okay :)

OpenStudy (anonymous):

\[\frac{p^2-4}{(p+2)^2}\]

hero (hero):

(p-2)/(p+2)

OpenStudy (anonymous):

factor the numerator and cancel. steps are \[\frac{9^2-4}{(p+2)(p+2)}=\frac{(p+2)(p-2)}{(p+2)(p+2)}=\frac{p-2}{p+2}\] what hero said!

OpenStudy (anonymous):

that 9 should be a 'p'

hero (hero):

Wow, somebody took my suggestion for "legend" congrats